If z="8(\cos \frac{5\pi}{3}+i\sin \frac{5\pi}{3})", then z6 is equal to


1
Expert's answer
2022-06-06T16:53:20-0400
"z^6=8^6\big(\cos(6(\dfrac{5\pi}{3}))+i\sin(6(\dfrac{5\pi}{3}))\big)"

"=262144\big(\cos(10\pi)+i\sin(10\pi)\big)=262144"